accompanying observations on x TOST time (hr) and y - RBOT time (min) for 12 oil specimens TOSTİ 4200 3600 3750 3650 4050 2795 RBOT 370 345 375 310 35 195 TOSTI 4870 4475 3450 2675 3750 3300 RBOT 400 375 285 220 345 280 (a) Calculate the value of the sample correlation coefficient. (Round your answer to four decimal places.)
media%2Ff83%2Ff83a329b-8a38-43b0-b5f5-a0
Answer #1

Solution: The data given in the question statement is as follows. Computing additional columns TOST RBOT 4200 3600 3750 3650Test Statistic: /n-2 V1-r2 t = 0.9186 × t 7.351 12-2 V1-0.91862 Degrees of Freedom = df = n-2 12-2-10 From t-distribution tab